The natural uncertainty in the measurement of speed of an electron in an atom is

estimated to be 2.2x10^4 m/s
1
Expert's answer
2019-09-13T10:16:14-0400

We can apply Heisenberg uncertainty principle to find the natural uncertainty in the measurement of the particle's position:


"∆x∆p\geq\frac{\hbar}{2},\\
\space\\
∆x\geq\frac{\hbar}{2m_e∆v}=\frac{1.055\cdot10^{-34}}{2\cdot[9.109\cdot10^{-31}]\cdot[2.2\cdot10^{-34}]},\\
\space\\
∆x\geq2.631\cdot10^{-9}\text{ m}."